Please help me on my homework on age problem on linear equation in one variable:...."Six years ago Carl was three times as old as John.Eight years hence,four times Carl's age will be five times John's age.Find their present ages.





























Answer by rothauserc(4718)

let x be Carl's age and y be John's age, then
(x-6) = 3*(y-6)
4*(x+8) = 5*(y+8)
first equation becomes
x-6 = 3y-18
x - 3y = -12
x = 3y -12
substitute for x in second equation
4*(3y-12+8) = 5y+40
12y-16 = 5y+40
7y = 56
y = 8, then use first equation
x-6 = 3*(8-6)
x-6 = 6
x = 12
therefore
Carl is 12 and John is 8